He stand with one foot on ceramic flooring and one foot on a wool carpet, making contact over an area of 75.0cm2 with each foot. Both the ceramic and the carpet are 5.00 cm thick and are 12.0ºC on their bottom sides. At what rate must heat transfer occurs from each foot to keep the top of the ceramic and carpet at 38.0ºC?


1
Expert's answer
2021-03-31T09:40:31-0400
"P=\frac{kA(T_h-T_c)}{d}\\
P_w=\frac{(0.035)(75\cdot10^{-4})(38-12)}{0.05}=0.1365 W\\
P_c=\frac{(5.077)(75\cdot10^{-4})(38-12)}{0.05}=19.8\ W"